OpenStudy (anonymous):

The incidence of cystic fibrosis, a recessive genetic disorder in the Caucasian population of United States, is 1 in every 2,500 individuals. Find the number of heterozygous carriers. (p + q = 1, p2 + 2pq + q2 = 1)

OpenStudy (shrutipande9):

p is homozygous dominant q is honozygous recessive pq is hetero as it is recessive....q=1/2500=0.0004 now p+q=1 p+0.0004=1 p=0.996 now u can find pq...which is heterozygoues carrier